Systems Staff Meeting Notes 2/9/98

Systems Staff Meeting Notes
February 9, 1998

Present: Rich Bennett, Bill Covey, Destry Fudge, Debra Harris, Winston Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota Whitfield

  1. Billís PC is still on the fritz - hence, no agenda.
  2. Jason Monahan departed last Wednesday, he will keep us apprised of his journeys. Denise Bennett has a line on a possible student replacement - if that doesnít work out, weíll conduct our own search.
  3. Intellikeys are being requested for our 4 entrances, we may end up with a reduced subset.
  4. Cage is still up in the outside hallway.
  5. Winston is Birthday Coordinator for this year - procedure will be that the last one to receive a birthday cake will make/bring the next personís cake. We will indulge on the Monday prior to the staff memberís birthday.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Scanner in AFA - floppy drive has been replaced, but it still wonít work. Lee will test it out with a functional floppy to see if the problem could be a power supply. It is getting a zip drive device conflict.
  2. Mail problems - growing pains - staff need to be set up to post to lists.
  3. Printer for Judiaca - 6 MP -is logged in and ready to go.
  4. Maps printer is also ready to go.
  5. Property should be coming to survey equipment on Wednesday, February 11.
  1. Education - Gresh has finished with his part - we will need to drag a hub over there.
  2. SE Tower is finished to the patch panel. Destry has the majority of Special Collections hooked up, but some are still on the old wiring.
  3. North Tower & Conference Room - one jack is ready.
  4. Grand Reading Room - may be used for walk-ups, as public access is being discussed.
  5. Phones - sharing is off. Will be pulling one phone jack to each location - to "keep costs down". Any modification will end up costing all over again.
  6. Other Tower - missing some jacks to go out. Martha wants the 3rd floor Processing area in action by Feb. 15.
  7. MSL - held up - Bill picked up the plans for Swamp II from Barbara Oliver. Implementation is still a long ways off.
  8. CISCO switches still have not been delivered.
  9. Bill is checking into a Lanier networked copier/printer/scanner - capable of producing 40 pages per minute - for the Directors Office. Pricing currently unavailable. Question of power problem - John Humphrey is working on it.
Software Projects Status
  1. Brian Burton visited Winston, said he has handed in his letter of resignation to MSL - will tentatively appear in Systems around February 19.
  2. Sharon is working on the OPS transfer for Personnel. She will try to demo it for Melita this week so she can get some feedback.
  3. Winston is working on ILL.
  4. Winston will demo OPS tracking tomorrow in the East Conference Room at 9:30 AM.
DP Coordinator
  1. Patron has checked out an I/B/E/S tape - authorization came via email. Debra will keep hard & email copies for verification.
  1. Suzy returns for two weeks. Both Bill & Suzy will be out Feb.23 - March 9.
  2. Spring Break begins March 9.

Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris